OpenStudy (anonymous):

the two cylinders are similar. What is the length of the altitude of the larger cylinder? radius 8;2 altitude ?;7

OpenStudy (saifoo.khan):

cross multiply.

OpenStudy (saifoo.khan):

8 : 2 x : 7 therefore: 2x = 7* 8 x = 56/2 x = 26

OpenStudy (anonymous):

its 28